OpenStudy (jkbo):

Please help A linear equation has an x-intercept of (5,0) and a y-intercept of (0,-8). Determine the slope of the line and write the equation in standard form.  A.m = -8/5 ; 8x - 5y = 40 B.m = -8/5 ; 8x + 5y = 40 C.m = 8/5 ; 8x + 5y = 40 D.m = 8/5 ; 8x - 5y = 40

OpenStudy (compassionate):

Hello, do you know the first step?

OpenStudy (jkbo):

not at all @Compassionate

OpenStudy (compassionate):

>A linear equation has an x-intercept of (5,0) and a y-intercept of (0,-8). Determine the slope of the line and write the equation in standard form.  We know its asking for a linear equation . we know x(5, 0) & y(0, -8) Our first challenge is: >Determine the slope Okay, the slope formula is: \[m = \frac{ y _{2}-y_{1} }{ x_{2} -x_{1} }\] Where: \[y_{2} = -8 \rightarrow y_{1} = 0\] So, we know that y2 = -8 and y1 = 0 Can you guess what x2 and x1 are?

OpenStudy (jkbo):

5 and 0 @Compassionate

OpenStudy (compassionate):

actually, x2 = 0 and x1 = 5 Is this making sense so far? Good. Okay, now take the points and plug them into the slope equation. \[m= \frac{ y2 - y1}{ x2 - x1 }\]

OpenStudy (compassionate):

If you're having trouble let me know.

OpenStudy (jkbo):

I got 8/5

OpenStudy (compassionate):

\[m = \frac{ -8 - 0 }{ 0 - 5 } = \frac{ -8 }{ -5 } = \frac{ 8 }{ 5 }\] Okay, so we just did the first part of the equation by finding the slope. > and write the equation in standard form.  Okay, so, basically, when you have two points and a slope, you can use the point-slope formula. \[y - y_{1} = m(x - x_{1})\] We also know our points are [(5, 0), (0, -8)] Where y1 = 0 Where x1 = 5 m = 8/5 So now, just plug all that into the point-slope equation, and solve.

OpenStudy (jkbo):

ok so the answer would be C. @Compassionate

OpenStudy (compassionate):

Correct.

OpenStudy (jkbo):

@Compassionate thank you